tell whether the system has one solution infinitely many solutions or no solution.
17=11y+12x
12x+11y=14

Answer by TutorDelphia(193)

17=11y+12x
12x+11y=14
This can be rewritten as:
11y+12x=17
11y+12x=14
Just from looking at this, we can see that these lines are parallel. Lets see the equations in slope intercept form
11y+12x=17
11y=-12x+17
y=-(12/11)x+17/11
11y+12x=14
11y=-12x+14
y=-(12/11)x+14/11
We can see that both equations have a slope of -12/11 which makes them parallel
as we can see here: